OpenStudy (anonymous):

Water is draining from a small cylindrical tank into a larger one below it. The small cylindrical tank has a radius of 4 feet and a height of 6 feet; the large cylindrical tank has a radius of 8 feet and a height of 16 feet. The small tank is initially full of water, and the water drains out at a rate of 12 cubic feet per second. Note: The volume of a cylinder is V = π(r^2)h. A. Find the volume Vs of the water remaining in the small tank as a function of time. B. How long does it take for the small tank to completely empty? C. Let z be the depth of the water in the large tank, which...

OpenStudy (anonymous):

...empty. Compute dz/dt. D. What fraction of the total amount of water is in the large tank at time t = 6?

OpenStudy (anonymous):

@dan815 look at the first comment, it continues the rest of the question :)

OpenStudy (michele_laino):

I think another way to solve the problem is to start from this differential equation: \[\Large \frac{{d{V_1}}}{{dt}} + \frac{{d{V_2}}}{{dt}} = 0\] and substituting the formula for each volume V, we get: \[\Large \frac{{d{h_2}}}{{dt}} = - {\left( {\frac{{{r_1}}}{{{r_2}}}} \right)^2}\frac{{d{h_1}}}{{dt}}\]

OpenStudy (dan815):

am i wrong to assume that water is just draining out of the small one at 12 cubic feet/sec

OpenStudy (trojanpoem):

You were right, volume as function of time is V(t) = pi (r)^2h - 12 t

OpenStudy (dan815):

oh dang

OpenStudy (trojanpoem):

V = 0 to get the time when it's empty .

OpenStudy (anonymous):

So because it is draining out of the little one at a rate of twelve, it is draining into the larger one at a rate of 12? and this rate is dv/dt or v'(t)?

OpenStudy (michele_laino):

so we have: \[\Large \frac{{d{h_1}}}{{dt}} = - \frac{{12}}{{\pi r_1^2}} = - \frac{{12}}{{16\pi }}\]

OpenStudy (trojanpoem):

at t = 6 get the amount of water from v(t) you will get the water in the small one and and minus it from the total volume of samll you will get the water in the big one.

OpenStudy (trojanpoem):

Yeah , samll -12 , big + 12

OpenStudy (anonymous):

so t=6 is the time it takes for the small one to completely empty?

OpenStudy (dan815):

-16*

OpenStudy (trojanpoem):

Nope, when the samll is empty V(t) = 0

OpenStudy (anonymous):

oh, oh never mind, I see that now

OpenStudy (trojanpoem):

@Michele_Laino , Your method is right, But the question specifies that it wants it with V(t)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

So then how does it work throwing depth into the mix for part C?

OpenStudy (dan815):

okay so you know how the volume is changing in the big one, what can you say about the height then

OpenStudy (trojanpoem):

I really don't understand it at all. ask Michele or dan.

OpenStudy (trojanpoem):

The height is the height of water

OpenStudy (anonymous):

Part C was supposed to read: C. Let z be the depth of the water in the large tank, which initially was empty. Compute dz/dt.

OpenStudy (dan815):

right, you know that the dv/dt = 12 cubic feet/sec

OpenStudy (trojanpoem):

You can easily find it from the Volume = 2 pi r^2 h ( get dev for time) 2 pi r^2 are constants h is the height of water

OpenStudy (trojanpoem):

dV/dt = 2pir^2 dh/dt solve for dh/dt

OpenStudy (dan815):

how much height is that for every 12 cubic feet?

OpenStudy (trojanpoem):

We want rate of change of height ?

OpenStudy (anonymous):

and for sure dh/dt=dz/dt??

OpenStudy (dan815):

the radius is 8 so 12 = pi * 8^2 * height height = ?

OpenStudy (trojanpoem):

what is dz/dt ?

OpenStudy (dan815):

that tells you how much height you have changed for every 12 cubic feet of volume which is dz/dt

OpenStudy (anonymous):

rate of change of depth? so it is dh/dt? and height would then be 12/(64pi)?

OpenStudy (trojanpoem):

Yeah, but dh/dt for small is different from dh/dt for big so we have to solve for each separately

OpenStudy (dan815):

yeah

OpenStudy (trojanpoem):

Yeah

OpenStudy (dan815):

dont they want you to just say the change in height for the big one

OpenStudy (anonymous):

So we just solved for the small tank? But we used a radius of 8, which is from the large tank?

OpenStudy (dan815):

if its the small one do the same thing, you know the volume change over time is 12 cubic feet /sec so 12 cubic feet = pi*4^2 *height height = 12/16pi this means the hight is changing by 12/16pi every second

OpenStudy (trojanpoem):

In the big tank dv/dt = 12 we solved already for the big one.

OpenStudy (trojanpoem):

@dan815 , at the small one is will be -12/16pi the height is decreasing

OpenStudy (dan815):

right you can say that

OpenStudy (anonymous):

but not dv/dt for the large one. dh/dt like the question asks?

OpenStudy (trojanpoem):

12 = pi (8)^2 dz/dt ( that's the one your question asks for)

OpenStudy (anonymous):

Oh! Thank you

OpenStudy (trojanpoem):

yw

OpenStudy (anonymous):

So, part D then addresses the fraction of the water that is in the tank at t=6... Is that v(6)/maximizedv(t)?

OpenStudy (trojanpoem):

yep , v(6)

OpenStudy (trojanpoem):

You mean the percentage of water ? at t = 6 ?

OpenStudy (anonymous):

But not just v(6), right? Because that would only be a number and I need the fraction? So I would have to divide v(6) by the maximized volume found by v(t), yes?

OpenStudy (trojanpoem):

Fraction = v(6)/ v(0) if I didn't misunderstand

OpenStudy (trojanpoem):

It will be like a/b from the small one is in the big one.

OpenStudy (anonymous):

That makes sense, except that t=0 is empty in the large tank? I understand that the water is only what fills up the small thank, though... I'm honestly really confused by this part

OpenStudy (trojanpoem):

If I didn't misunderstand the question for ex the question wants how much from the water is currently is in the big like 2/3 or 1/3 so we current amount/ total but how to get the total ? put t = 0 when no water has drained from the small one.

OpenStudy (trojanpoem):

asks for*

OpenStudy (dan815):

why are you doing v(6)/v(0)

OpenStudy (trojanpoem):

Current water / total water hmm ?

OpenStudy (dan815):

v(0) for the big tank is = remember

OpenStudy (dan815):

v(0) for big tank = 0

OpenStudy (trojanpoem):

I am doing v(0) for the samll tank

OpenStudy (trojanpoem):

v(t) = pi (4)^2 - 12 t

OpenStudy (trojanpoem):

It's like current water/ volume of small

OpenStudy (dan815):

i thought D was asking for large tank

OpenStudy (trojanpoem):

It's asking for the fraction of water in the large so we have to get the total water and the current in the big am I mistaken ?

OpenStudy (anonymous):

It is asking for large tank so would it look like: pi (8)^2 + 12 (6) /pi (4)^2 - 12 (0) ?

OpenStudy (anonymous):

Which was my hope of putting v(t) into large tank terms for v(6) and dividing it by v(t) for the small tank at v(0)? hopefully?

OpenStudy (trojanpoem):

I was thinking of get 12 * 6/ 4^2 pi

OpenStudy (anonymous):

why?

OpenStudy (trojanpoem):

12 * 6 ( amount lost) / total volume

OpenStudy (trojanpoem):

@dan815 , come.

OpenStudy (dan815):

ah i gotcha!

OpenStudy (trojanpoem):

pi (8)^2 + 12t < means we are adding more than the volume of the large

OpenStudy (trojanpoem):

Check with me number D dan

OpenStudy (dan815):

you are doing it right

OpenStudy (dan815):

i have bad habit of not reading question lol, i just read what fraction and stopped lol, and assumed they wanted the fraction of the height of the big tank

OpenStudy (dan815):

its called selective reading xD u only read words here and there, u tend to do it when u gotta read so many questions over and over

OpenStudy (trojanpoem):

xD

OpenStudy (trojanpoem):

so they want fraction in the big , find we get amount in big / volume of big

OpenStudy (anonymous):

But they only want the fraction of the water, which would be amount in big/ volume of small

OpenStudy (dan815):

yes water

OpenStudy (trojanpoem):

Ey selective reading fails dude -_-

OpenStudy (dan815):

so do v(t)= pi*4^2*6-12t 1- (V(6)/V(0))

OpenStudy (dan815):

V(6)/V(0) gives you the fraction in the small tank so 1- that is the fraction in the big tank

OpenStudy (trojanpoem):

I agree with dan815 ++

OpenStudy (dan815):

gimme medal + monay plz

OpenStudy (anonymous):

what I'm not understanding is that it can't be the same v(t) because one is referring to the big tank and one is referring to the small tank? I'm confused as to how the 1- fixes that?

OpenStudy (trojanpoem):

Let's say v(6)/ v(0) is 2/3 of the volume in the water so in the big is 1/3

OpenStudy (anonymous):

Oh! Gotcha. Thank you both so much!

OpenStudy (trojanpoem):

yw

OpenStudy (dan815):

sure thing :)